# MASS: Motion-Aware Spatial–temporal Grounding for Physics Reasoning and Comprehension in Vision-Language Models

### v1

@@ 4.2 Spatial Motion Feature Extraction
Leveraging the narrowed-down, visually grounded entities identified by the probing questions, we perform motion tracking for each entity within its video sub-sequences using CoTracker3 [karaev2024cotracker3], which tracks arbitrary point trajectories inside the grounded regions. For each entity, we compute the spatial deviations across time, including averaged motion vectors and the first/last tracked positions. In parallel, we estimate per-frame depth using Depth Anything V2 [depth_anything_v2] to provide spatial awareness. By aligning sampled RGB frames with their corresponding depth maps, we project the tracked motion vectors over each tracking points throughout the entity into 3D space, yielding both 3D positions and full spatio-temporal motion trajectories for each entity.
@@ 4.3 Visual Feature Representation
After encoding visual features across different modules, the key challenge lies in integrating heterogeneous spatial–motion outputs into the VLM pipeline without disrupting its native architecture. Prior work [wu2024autohallusion, li2025videohallu] shows that LLM backbones better comprehend abstract, factual, and physics-related concepts when such information is expressed in the language space rather than as raw visual features.
Motivated by this, we fuse visual cues through natural-language representations aligned with the tokenized questions. We compute these features because VLMs cannot reliably infer fine-grained spatial layouts or motion dynamics from compressed visual embeddings; providing explicit, structured cues supplies the physical information needed for stable and interpretable reasoning. Our visual-feature serialization converts spatial–motion signals into structured motion-trajectory sequences for each entity, including 3D start/end positions, motion vectors, bounding boxes within each video sub-segment, and associated temporal indices. These attributes are expressed through predefined natural-language templates, enabling the VLM to access precise spatial–temporal information in the modality it handles most effectively. We provide the visual feature representation template incorporated into the question–answering format used for VLM training in Appendix C.
@@ 4.4 Training Pipeline
Although visual grounding and spatial–motion features provide essential cues, VLMs cannot directly leverage these signals for higher-level reasoning because their representations are misaligned with those learned during pre-training. As a result, even precise motion trajectories, spatial positions, and temporal indices—such as knowing an object is intact at position A at time 1 and broken at position B at time 2—cannot be meaningfully synthesized without additional alignment. The model must bridge this representational gap to integrate structured cues, filter irrelevant details, and reason over scene evolution rather than merely retrieving or describing visual content.
To address this gap, we perform post-training to enhance the VLM’s reasoning capabilities. Chain-of-thought (CoT) reasoning [wei2022chain] helps the model process complex temporal and physical concepts more effectively. By encouraging intermediate reasoning steps, CoT aligns the VLM’s decision process with the multi-stage deductions needed for physical comprehension. This training enables the VLM to operate on grounded spatial–motion features, integrate temporal changes, and produce coherent, physics-aware explanations rather than treating cues as isolated descriptors. We evaluate both Supervised Fine-tuning (SFT) and Reinforcement Fine-tuning (RFT). For RFT, we adopt Temporal Group Relative Policy Optimization (T-GRPO) [feng2025video], which adds temporal-aware rewards on top of vanilla GRPO [shao2024deepseekmath], a key component for video understanding. We further incorporate ROUGE [lin2004rouge] for semantic coherence, a temporal reward to strengthen temporal reasoning, and a format reward to encourage proper CoT generation. Implementation details of our method are provided in Appendix B.
@@ 5 Empirical Results
For our experiments, we evaluate $14$ widely used state-of-the-art VLMs as baselines (e.g., GPT-4o and Gemini-2.5-Flash), covering 7B and 13B parameter scales. We report performance comparisons between these baselines and VLMs augmented with MASS in Table 2, and present ablations on our backbone models and post-training strategies in Table 3. For evaluation, we use GPT-4.1-Mini as an automated judge to assess the factual correctness between each model’s answer and the ground-truth annotations regarding these physical comprehension and reasoning questions provided. We provide the evaluation template in Appendix D. We provide additional qualitative analyses of question–answering results and extended evaluations in the Appendix. Appendix E reports results on general-purpose real-world video QA benchmarks, while Appendix F includes qualitative video question–answering examples.
Spatial–Temporal Signals Are Critical for Physics Understanding. We evaluate baseline VLMs in the 7B–8B range, including LLaVA [liu2023visual], Qwen-VL [bai2025qwen2], and InternVL [zhu2025internvl3], with additional comparisons to larger InternVL3.5 models. Surprisingly, newer or larger variants (e.g., Qwen2.5-VL-7B vs. Qwen3-VL-8B, or InternVL3-8B vs. InternVL3.5-8B-Flash) do not consistently yield better performance under identical settings. Scaling further offers limited benefit: within the InternVL3.5 family, increasing from 8B to 38B yields only a modest $4.0\%$ gain, and Qwen3-VL-32B performs similarly to its 8B counterpart. These results indicate that model scale alone does not reliably strengthen physics reasoning. In contrast, integrating MASS produces substantial improvements. Qwen2.5-VL-7B and LLaVA-OneVision-7B outperform the previous best model (Qwen3-VL-8B) by $8.7\%$ and $6.0\%$, respectively. Notably, Qwen2.5-VL-7B+MASS matches Gemini-2.5-Flash and surpasses it on physics abnormality detection—the most reasoning-intensive category. These findings demonstrate that explicit spatial–temporal representations provide the structured physical cues VLMs inherently lack, enabling more faithful physics-driven reasoning.
@@ 5.1 Baseline Insights: Why Motion-Aware Spatial-temporal Grounding Matters
Table 2: Accuracy (%) across categories in MASS-Bench: Performance of VLMs across five physics-related categories—Spatial Understanding (SU), Temporal Understanding (TU), Motion & Action Recognition (MAR), Physics Comprehension (PC), and Physical Abnormality Detection (PA). Models are grouped by family and sorted by overall accuracy. Integrating MASS with Qwen2.5-VL-7B and LLaVA-OneVision-7B substantially improves performance over their baselines and achieves accuracy comparable to Gemini-2.5-Flash, the current SoTA in video understanding, by leveraging explicit spatial–temporal representations that enhance physics comprehension and reasoning.
@@ 5.2 Ablation Study: The Role of Reasoning
Reasoning Dominates Beyond the Base Model. Table 3 compares post-training results on two backbone VLMs, Qwen2.5-VL-7B [bai2025qwen2] and LLaVA-OneVision-7B [li2024llava]. Across both models, enhanced reasoning contributes far more than the backbone choice itself: pairing Qwen2.5-VL-7B with MASS yields a $9.6\%$ improvement, while LLaVA-OneVision-7B combined with MASS achieves a larger $13.5\%$ gain. After post-training, the base model’s influence becomes secondary, as physics comprehension, especially challenging tasks like physics abnormality detection, demands integrated reasoning over visual physics cues, spatial–temporal structure, and commonsense priors. This synergy, rather than raw model capacity, ultimately drives the performance improvements.
Why RFT Matters: SFT Alone Is Not Enough. Table 3 compares SFT and RFT strategies. While RFT (via GRPO) consistently boosts performance, SFT leads to notable degradation: $8.6\%$ on Qwen2.5-VL-7B and $2.9\%$ on LLaVA-OneVision-7B. This highlights that simple supervised alignment cannot equip VLMs to handle complex physics reasoning. Physics comprehension, especially detecting abnormal or non-intuitive dynamics, requires models to integrate grounded visual cues with spatial–temporal structure and commonsense priors. RFT explicitly strengthens this reasoning process, whereas SFT merely memorizes correlations between input–output pairs, limiting its effectiveness on physics-intensive tasks.
@@
Table 3: Ablation of Post-Training Strategies. Performance comparison of base VLMs, SFT variants, and GRPO-enhanced models across five physics-related categories. GRPO consistently provides substantial gains over both the base models and their SFT counterparts, whereas SFT generally degrades performance.
@@ 5.3 Discussion
Challenges in Physical Abnormality Detection Remain. A key observation from Table 2 and Table 3 is that Qwen2.5-VL-7B and LLaVA-OneVision-7B integrated with MASS outperform all other VLMs, including Gemini-2.5-Flash, on physical abnormality detection. However, compared with more grounded physics comprehension tasks, abnormality detection is inherently more difficult. It requires VLMs to overcome hallucinations rooted in their learned priors about real-world physics [li2025videohallu] and to reason about why a visual process violates physical laws. While our method yields a notable $12\%$ improvement on this task, indicating the promise of motion-aware spatiotemporal grounding for enhancing perception and reasoning, further advances are needed to robustly handle these challenging cases.
@@ 6 Conclusion, Limitations and Future Work
We present a model-agnostic framework that bridges raw video content and physics-aware reasoning by converting physical cues into structured, interpretable inputs suitable for VLM perception and inference. To support this, we introduce a comprehensive benchmark of physics-oriented video QA pairs with fine-grained annotations, including visual detections, sub-segment grounding, and full-sequence 3D motion tracking. To our knowledge, MASS-Bench is the first dataset providing both positive and negative demonstrations paired with questions on spatial–temporal understanding, physics reasoning, and abnormality detection. Our method, MASS, injects spatial–temporal signals into the VLM language space through depth-based 3D encoding, visual grounding, and motion-aware trajectory modeling, yielding consistent gains over strong baselines and recent SoTA VLMs on real-world and AIGC videos. Our approach has limitations. Although post-training improves physics reasoning, models still struggle with comprehensive real-world dynamics and commonsense cues. The current implementation also faces challenges in crowded scenes, where multi-object tracking increases computational overhead. In addition, subtle cues in reasoning-heavy cases may be overshadowed by language priors, leading to hallucinations. Future work includes enhancing long-range motion grounding and temporal reasoning to better capture extended physical dynamics, improving multi-object tracking via more scalable spatial–temporal profiling, and expanding high-quality training data with diverse positive/negative examples. These efforts aim to further reduce hallucinations, improve physical fidelity, and broaden applicability to complex physics-centered video reasoning.
[…]
@@ Appendix A Additional Details of MASS-Bench
This section supplements the dataset description in Section 3 by providing expanded definitions and illustrative examples for each video question-answering pairs.
A.1 Dataset Composition
MASS-Bench aggregates diverse physics-centric video sources to ensure broad coverage across spatial, temporal, and motion-driven phenomena. In total, the dataset comprises 1,2291{,}229 samples ($14.63\%$) from MotionSight [du2025motionsight], 4,0094{,}009 samples ($47.95\%$) from VideoPhy2 [bansal2025videophy], and 2,8622{,}862 samples ($34.23\%$) from VideoHallu [li2025videohallu].
A.2 Example Type Composition
To characterize the types of physics reasoning required in MASS-Bench, we follow the definitions introduced in Section 3 and categorize each video question-answer pair into one of two groups.
Positive Examples capture scenarios in which the physical dynamics depicted in the video are consistent with real-world physics. These include both real and synthetically generated videos whose motion patterns follow physically plausible trajectories. Questions in this category primarily evaluate a model’s ability to comprehend and reason about correct physical processes. MASS-Bench contains a total of 3,4363{,}436 such examples ($41.10\%$).
Negative Examples correspond to cases where at least one aspect of the video exhibits a physical abnormality, such as trajectories or interactions that violate real-world physics. Successfully answering these questions requires VLMs to overcome language priors and reason directly from observed visual evidence. This category includes 4,9254{,}925 examples ($58.90\%$).
Overall, the two groups jointly evaluate physics comprehension under both physically valid and physically inconsistent conditions.
A.3 Question Type Composition
To assess the range of reasoning skills required by VLMs, we categorize all video–question–answer pairs in MASS-Bench into two types. Representative examples of each are shown in Figure 2.
Factual Questions focus on concrete, visually grounded information related to spatial layout, temporal ordering, object motion, or physical abnormalities. Examples include queries such as “What is the spatial position of the watermelon in the video?” or “How does the watermelon move after it is shot?” These questions typically contain explicit cues that narrow the perceptual scope, enabling the model to identify the relevant entities and perform targeted grounding and reasoning. This category consists of 5,4275{,}427 examples ($67.0\%$).
Critical Thinking Questions require higher-level inference and more abstract reasoning. They often omit explicit grounding cues and instead ask broadly scoped questions such as “What is the watermelon doing in this video?” or “What is abnormal in the scene?” Successfully answering them requires the model to infer intent, identify salient events, and interpret physical dynamics without direct guidance from the query. This category contains 2,6732{,}673 examples ($33.0\%$).
Together, these two question types enable a comprehensive evaluation of both grounded perception and higher-order physics reasoning in VLMs.

### v2

@@ 4.2 Spatial Motion Feature Extraction
Leveraging the narrowed-down, visually grounded entities identified by the probing questions, we perform motion tracking for each entity within its video sub-sequences using CoTracker3 [25], which tracks arbitrary point trajectories inside the grounded regions. For each entity, we compute the spatial deviations across time, including averaged motion vectors and the first/last tracked positions. In parallel, we estimate per-frame depth using Depth Anything V2 [62] to provide spatial awareness. By aligning sampled RGB frames with their corresponding depth maps, we project the tracked motion vectors over each tracking points throughout the entity into 3D space, yielding both 3D positions and full spatio-temporal motion trajectories for each entity.
@@ 4.3 Visual Feature Representation
After encoding visual features across different modules, the key challenge lies in integrating heterogeneous spatial–motion outputs into the VLM pipeline without disrupting its native architecture. Prior work [57, 36] shows that LLM backbones better comprehend abstract, factual, and physics-related concepts when such information is expressed in the language space rather than as raw visual features.
Motivated by this, we fuse visual cues through natural-language representations aligned with the tokenized questions. We compute these features because VLMs cannot reliably infer fine-grained spatial layouts or motion dynamics from compressed visual embeddings; providing explicit, structured cues supplies the physical information needed for stable and interpretable reasoning. Our visual-feature serialization converts spatial–motion signals into structured motion-trajectory sequences for each entity, including 3D start/end positions, motion vectors, bounding boxes within each video sub-segment, and associated temporal indices. These attributes are expressed through predefined natural-language templates (detailed in the supplementary material), enabling the VLM to access precise spatial–temporal information in the modality it handles most effectively.
@@ 4.4 Training Pipeline
Although visual grounding and spatial–motion features provide essential cues, VLMs cannot directly leverage these signals for higher-level reasoning because their representations are misaligned with those learned during pre-training. As a result, even precise motion trajectories, spatial positions, and temporal indices such as knowing an object is intact at position A at time 1 and broken at position B at time 2 cannot be meaningfully synthesized without additional alignment. The model must bridge this representational gap to integrate structured cues, filter irrelevant details, and reason over scene evolution rather than merely retrieving or describing visual content.
We perform post-training to enhance the VLM’s reasoning capabilities to bridge this gap. Chain-of-thought (CoT) reasoning [55] helps the model process complex temporal and physical concepts more effectively. By encouraging intermediate reasoning steps, CoT aligns the VLM’s decision process with the multi-stage deductions needed for physical comprehension. This training enables the VLM to operate on grounded spatial–motion features, integrate temporal changes, and produce coherent, physics-aware explanations rather than treating cues as isolated descriptors. We evaluate both Supervised Fine-tuning (SFT) and Reinforcement Fine-tuning (RFT). For RFT, we adopt Temporal Group Relative Policy Optimization (T-GRPO) [20], which adds temporal-aware rewards on top of vanilla GRPO [48], a key component for video understanding. We further incorporate ROUGE [39] for semantic coherence, a temporal reward to strengthen temporal reasoning, and a format reward to encourage proper CoT generation.
@@ 5 Empirical Results
For our experiments, we evaluate 12 widely used state-of-the-art VLMs as baselines (e.g., GPT-4o and Gemini-2.5-Flash), covering 7B–38B parameter scales. We report performance comparisons between these baselines and VLMs augmented with MASS in Table 1, and present ablations on our backbone models and post-training strategies in Table 3. For evaluation, we use GPT-4.1-Mini as an automated judge to assess the factual correctness between each model’s answer and the ground-truth annotations regarding these physical comprehension and reasoning questions provided. We provide additional ablation studies are provided in the supplementary material, including real-world video QA generalization (Section E), noise-induced robustness analysis (Section F), and inference efficiency measurements (Section G).
@@
Table 1: Accuracy (%) across categories in MASS-Bench: Performance of VLMs across five physics-related categories—Spatial Understanding (SU), Temporal Understanding (TU), Motion & Action Recognition (MAR), Physics Comprehension (PC), and Physical Abnormality Detection (PA). Models are grouped by family and sorted by overall accuracy. Scaling to 32B–38B parameters yields only marginal gains over 8B counterparts, whereas integrating MASS with 7B models substantially improves performance and achieves accuracy comparable to Gemini-2.5-Flash by leveraging explicit spatial–temporal representations.
Spatial–Temporal Signals Are Critical for Physics Understanding. We evaluate baseline VLMs across a wide parameter range, from 7B–8B models in the LLaVA [40], Qwen-VL [3], and InternVL [70] families, up to 32B–38B variants and close-source models such as GPT-4o and Gemini-2.5-Flash. Under the same evaluation setting, neither newer variants nor larger models consistently perform better. For instance, Qwen2.5-VL-7B and Qwen3-VL-8B achieve similar results, and InternVL3.5-8B-Flash shows no clear advantage over InternVL3-8B. This trend also appears at larger scales: InternVL3.5-38B-Flash improves by only $4.0\%$ over its 8B version, while Qwen3-VL-32B remains comparable to Qwen3-VL-8B. By contrast, incorporating MASS yields substantial improvements. Qwen2.5-VL-7B+MASS and LLaVA-OneVision-7B+MASS exceed the previous best model, Qwen3-VL-8B, by $8.7\%$ and $6.0\%$, respectively. Qwen2.5-VL-7B+MASS also matches Gemini-2.5-Flash overall and outperforms it on physics abnormality detection, the most reasoning-heavy category. Together, these results show that explicit spatial–temporal representations supply physical cues that standard VLMs often miss, and are therefore key to reliable physics reasoning.
@@ 5.2 Ablation Study: The Role of Reasoning
Reasoning Dominates Beyond the Base Model. Table 3 compares post-training results on two backbone VLMs, Qwen2.5-VL-7B [3] and LLaVA-OneVision-7B [26]. Across both models, enhanced reasoning contributes far more than the backbone choice itself: pairing Qwen2.5-VL-7B with MASS yields a $9.6\%$ improvement, while LLaVA-OneVision-7B combined with MASS achieves a larger $13.5\%$ gain. After post-training, the base model’s influence becomes secondary, as physics comprehension, especially challenging tasks like physics abnormality detection, demands integrated reasoning over visual physics cues, spatial–temporal structure, and commonsense priors. This synergy, rather than raw model capacity, ultimately drives the performance improvements.
Visual Evidence and Motion Context Are Complementary. We ablate the input modality using three settings: (i) Text-only (motion profile + question), (ii) Visual-only (video tokens + question), and (iii) Both (video tokens + motion profile + question, our default). Table 2 reports these results alongside the SFT vs. GRPO comparison. Text-only consistently underperforms visual-only across all models (e.g., GPT-5.2: 42.09% vs. 56.17%; Gemini-2.5-Pro: 43.39% vs. 63.41%), confirming that many questions are inherently video-dependent. While the motion profile provides useful temporal summaries, it cannot replace the spatial detail and grounded evidence in the original video. Consistent with this, about 26% of text-only outputs are judged as uncertain by the LLM judge. Using both modalities yields the best results. Under the Both setting, GRPO + MASS improves Qwen2.5-VL-7B by 9.6% and LLaVA-OneVision-7B by 13.0%. This shows that the motion profile is most effective as complementary context rather than a substitute for video. The benefit also depends on model capacity: large models such as GPT-5.2 and Gemini-2.5-Pro can exploit motion-profile text even without post-training, whereas 7B models gain little from the added text unless they are further trained to integrate it. This pattern suggests that the improvement does not come from extra textual context alone, but from learning to align motion-profile cues with visual evidence for better physics reasoning.
Why RFT Matters: SFT Alone Is Not Enough. Table 3 compares SFT and RFT strategies. SFT yields limited benefit overall: it slightly reduces performance for Qwen2.5-VL-7B by 0.3% and brings only a modest 4.2% gain for LLaVA-OneVision-7B, leaving both close to the off-the-shelf models. By contrast, GRPO produces large and consistent gains across both backbones, reaching 61.93% on Qwen2.5-VL-7B and 59.24% on LLaVA-OneVision-7B. We believe this is because GRPO optimizes task-level rewards directly, making it better suited than SFT for learning physical reasoning that depends on temporal evidence and causal consistency. The gains are especially clear on SU, TU, MAR, and PC, where stronger reasoning is needed to connect motion cues with physical dynamics. PA is the main exception, where SFT performs better in our reruns, likely because PA depends more on answer calibration: the model must not only detect the abnormality, but also describe it in a precise free-form way that matches the reference. Even so, GRPO remains the stronger strategy overall.
Input Variants Training Strategy
Model Text Visual Both SFT GRPO
GPT-5.2 42.09 56.17 59.46
Gemini-2.5-Pro 43.39 63.41 64.30
Qwen2.5-VL-7B 32.14 52.30 52.38 50.40 52.89
+ GRPO + MASS 39.24 52.94 61.93 52.01 61.93
LLaVA-OneVision-7B 31.69 45.79 46.21 47.60 50.33
+ GRPO + MASS 35.63 46.72 59.24 50.03 59.24
Table 2: Input modality ablation (left) and training strategy comparison (right) for models with and without MASS. Results are reported as accuracy (%) on MASS-Bench.
@@
Table 3: Ablation of post-training strategies. Accuracy (%) of base VLMs, SFT variants, and GRPO-enhanced models across five physics-related categories. GRPO provides substantial gains over both base models and SFT in most categories.
Training Data and Module Ablations. We ablate the GRPO training data and the motion representation under the same training budget. Table 4 shows two main takeaways. First, the balanced Both setting works best on both backbones. Real-only training brings almost no gain over the off-the-shelf models, while Syn-only helps more but still remains below Both. This suggests that the two data sources help in different ways: real videos add natural variation in appearance and motion, while synthetic data more clearly expose the physical patterns the model needs to learn. Second, 3D motion matters. Replacing the full 3D motion representation with 2D motion lowers accuracy for both models. Because this ablation only removes the depth-aware projection, the drop is not just about having fewer useful tokens; it shows that depth-aware motion provides information that 2D motion alone misses. Overall, the gains come from combining complementary training data with motion cues that better reflect the physical structure of the scene.
Baselines (MASS) Training Data Modules
Model (+ MASS) w/ w/o Real-only Syn-only 2D Motion
Qwen2.5-VL-7B 61.93 52.30 52.40 56.30 56.50
LLaVA-OneVision-7B 59.24 45.79 49.92 53.41 52.30
Table 4: Ablation of GRPO training data and depth module. Accuracy (%) on MASS-Bench. w/ and w/o denote the full MASS pipeline (real+syn data, 3D motion) and the base model without MASS, respectively. Real-only and Syn-only vary the training data, while 2D Motion removes the depth-aware 3D projection.
@@ 5.3 Discussion
Challenges in Physical Abnormality Detection Remain. A key observation from Table 1 and Table 3 is that Qwen2.5-VL-7B and LLaVA-OneVision-7B integrated with MASS outperform all other VLMs, including Gemini-2.5-Flash, on physical abnormality detection. However, compared with more grounded physics comprehension tasks, abnormality detection is inherently more difficult. It requires VLMs to overcome hallucinations rooted in their learned priors about real-world physics [36] and to reason about why a visual process violates physical laws. While our method yields a notable $12\%$ improvement on this task, indicating the promise of motion-aware spatiotemporal grounding for enhancing perception and reasoning, further advances are needed to robustly handle these challenging cases.
@@ 6 Conclusion, Limitations and Future Work
We present a model-agnostic approach that bridges the gap between raw video content and physics-aware reasoning by translating physical-world cues into structured representations compatible with VLM inference. We introduce MASS-Bench, to our knowledge the first benchmark pairing positive and negative video demonstrations with QA on spatial-temporal comprehension, physics-based reasoning, and abnormality detection, supported by fine-grained visual detections, sub-segment grounding, and 3D motion tracking. Our method, MASS, injects spatial–temporal signals into the VLM language space via depth-based 3D encoding, visual grounding, and motion-aware trajectory modeling, yielding consistent gains over strong baselines and recent SoTA VLMs on both real-world and AIGC videos.
Limitations. Although post-training improves reasoning, it does not fully solve physical understanding in real-world settings. Performance may still decline in crowded scenes, where multi-object tracking becomes more demanding, and hallucinations persist in reasoning-heavy cases when subtle visual evidence is misread or overshadowed by language priors. MASS-Bench is entity-centric, focusing on the top-$K$ targets, and thus does not evaluate highly crowded scenes. Robust handling of strong camera ego-motion also remains open. Encouragingly, MASS remains reasonably stable under noisy motion cues, with gradual rather than catastrophic degradation (see Section F in Supplementary Material).
Future Work. Future work will improve long-range motion grounding and temporal reasoning in longer videos, and make multi-object tracking more scalable in crowded scenes through better spatial–temporal profiling. We also plan to extend the benchmark to denser multi-entity settings and stronger camera motion. Finally, expanding the training data with more diverse positive and negative examples may further reduce hallucinations and improve coverage of more challenging physics understanding tasks.
@@ Appendix A Additional Details of MASS-Bench
This section supplements the dataset description in Section 3 with expanded definitions and statistics.
A.1 Dataset Composition and Metadata
MASS-Bench aggregates three physics-centric video sources: 1,2291{,}229 samples ($14.7\%$) from MotionSight [17], 4,0094{,}009 ($47.9\%$) from VideoPhy2 [5], and 2,8622{,}862 ($34.2\%$) from VideoHallu [36]. In total, the dataset comprises 6,0936{,}093 training and 2,2682{,}268 test examples drawn from 4,3504{,}350 unique videos. Videos average $545.8$ frames (${\sim}19.6$ s at $27.4$ FPS) with a mean resolution of $1120{\times}702$ pixels. Due to overlap inherited from the source datasets, approximately $1.5\%$ of unique videos appear in both splits; however, no (video, question) pairs are shared across splits, so there is no direct QA leakage. All methods are evaluated on the same fixed test set with identical evaluation scripts. Importantly, MASS relies only on depth and tracking signals extracted from the input video at inference time and does not use any ground-truth test annotations.
A.2 Question Categories
We group all questions into five categories reflecting the type of physical reasoning required, ordered from basic perception to advanced comprehension. Table 5 shows the distribution.
Category Count Pct.
SU (Spatial Understanding) 2,785 33.3%
TU (Temporal Understanding) 1,633 19.5%
PA (Physical Abnormality Detection) 1,432 17.1%
PC (Physics Comprehension) 1,304 15.6%
MAR (Motion & Action Recognition) 1,205 14.4%
Table 5: Distribution of question categories in MASS-Bench.
These categories are ordered from easy to challenging. VLMs must first establish spatial and temporal awareness, then recognize motion patterns, before progressing to physics comprehension and violation detection.
A.3 Example and Question Types
We categorize each video–question–answer pair along two complementary axes.
For example type, we distinguish positive examples (3,4363{,}436, $41.1\%$), where the video follows real-world physics, from negative examples (4,9254{,}925, $58.9\%$), where at least one part of the scene violates physical plausibility. Negative examples are more challenging because the model must rely on visual evidence rather than defaulting to language priors.
For question type (Figure 2), we separate factual questions (5,4275{,}427, $67.0\%$) from critical-thinking questions (2,6732{,}673, $33.0\%$). Factual questions provide explicit cues that focus attention on particular entities or events. Critical-thinking questions are less direct: they omit such cues and instead require the model to infer intent, identify salient events, or reason about the underlying physical dynamics.
A.4 Motion Grounding Annotations
Beyond the question, video, and ground-truth answer, each pair includes five layers of motion-grounding annotation: (1) temporal video segmentation indexed by frame ranges; (2) visual grounding for each queried entity, specified by entity name and bounding box; (3) an entity-level temporal profile that tracks each grounded entity across the full video, persisting once detected and left blank in segments where the entity is absent; (4) per-segment motion attributes, including the first and last observed 3D positions to capture coarse spatial layout; and (5) 3D motion vectors representing each entity’s temporal displacement. Together, these cues convert physics-intensive perceptual challenges into structured textual and mathematical representations that support more reliable physical reasoning.